HUBBLE SPACE TELESCOPE DAILY REPORT #2211 PERIOD COVERED: 0000Z (UTC) 09/17/98 - 0000Z (UTC) 09/18/98 Daily Status Report as of 261/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Two Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed Two Sets of NIC/2 7894 (Completing A Near-Infrared Search for Very Low Mass Companions to Stars within 10 pc) The Near Infrared Camera (Camera Two) was used to make observations searching for very low mass (VLM) companions to the star CD-32D1329. This search will be the largest, most sensitive, volume-limited search for VLM companions ever undertaken. The observations were completed as planned, and no anomalies were reported. 1.4 Completed WF/PC-2 7386 (A Snapshot Survey of Nova Shells) The WF/PC-2 was used to make observational images of the super nova shell around variable star QU-Val in order to detect and measure its expanding gaseous shell. These observations will trace the distribution of hot and cool gas in the nova shells, and show their degree of inhomogeneity on both large and small scales. The observations were completed as scheduled, and no problems were noted. 1.5 Completed FGS 7491 (The Masses of the Brown Dwarf Candidate GL 623B) The FGSs were used to conduct astrometry at 1-mas accuracy of faint companions to nearby M-dwarfs. This is the most productive direct method to determine accurate masses of stars near the end of the main sequence. From preliminary FGS results we estimate the mass of GL 623B as 0.10 solar masses, placing both near the stellar/substellar border. Such a mass determination for GL 623B should establish conclusively its nature as star or substellar object. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed Three Sets of STIS/CCD 8062 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude observations take in parallel with the prime science observations. The observations were completed as planned, and no problems were reported. 1.7 Completed Two Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.8 Completed Two Sets of STIS/CCD 8064 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal - Backup) The Space Telescope Imaging Spectrograph (CCD) was used to conduct a non-scripted parallel POMS test. The observations were completed as scheduled, and no problems were reported. 1.9 Completed NIC/1/2/3 7245 (Galaxy Evolution in Rich Clusters) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the galactic cluster CL0024+16. The purpose is to study the morphological and spectral evolution of galaxies within this cluster environment. A dark calibration image was also taken. The observations were completed as planned, and no problems were reported. 1.10 Completed WF/PC-2 6603 (Probing Proto-Planetary Disks in the Orion Nebula) The WF/PC-2 was used to make images of the proto-planetary disk NGC-1976 which is located in the Orion Nebula. This is to study the fundamental physical properties and structures of recently discovered, and uniquely visible, proto-planetary disks around young stars. The observations were completed as planned, and no problems were reported. 1.11 Completed Two Sets of NIC/2/3 8082 (NICMOS pure parallel Fall 98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel observations while another instrument is prime. The observations were executed as scheduled, and no anomalies were noted. 1.12 Completed NIC/1/2/3 7233 (Dust Disks Around Main Sequence Stars) The Near Infrared Camera (Camera One, Two and Three) was used to make observations of the dust disk around the main sequence star HD34700. The observations were made with the coronagraph to minimize the effects of glare from the bright central star. A dark calibration image was also taken. The observations were completed as scheduled, and no anomalies were reported. 1.13 Completed NIC/1/2/3 7899 (An IR Search For faint Companions to Pleiades Very Low-Mass Stars and Brown Dwarfs) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of MH01. This is part of a infrared campaign to search for faint companions to Pleiades very low-mass stars and brown dwarfs. The Pleiades has been the best site for brown dwarf {BD} searches, with half of the presently "confirmed" BDs being Pleiades members. The observations were executed as scheduled, and no anomalies were noted. 1.14 Completed STIS/CCD 7579 (Coronographic Imaging & Spectroscopy of Beta Pictoris) The Space Telescope Imaging Spectrograph (CCD) was used to make coronographic imaging and spectrographic analysis of the star HD39060 (a.k.a., Beta Pictoris). The program aims to obtain STIS/G430L spectra of the central region of the Beta Pic disk. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 5 Successful : 5 Scheduled Reacquisitions : 10 Successful : 10 2.2 FHST Updates: Scheduled : 10 Successful : 10 2.3 Operations Notes: The SSR EDAC error counter was reset three times using ROP SR-1. An MA/GN mode SSR dump test was successfully supported at 260/1615Z. PACOR-II reported that data quality was nominal. Again, the HST IP was used as the prime system from approximately 260/1200Z until approximately 260/2000Z. There were no significant problems. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH